Sunday 01/13/13

3:30P – 11:55 P

Fly work

  • Flipped most of rescue crosses.   Several vials extinct.  Should still have more vials than I want to screen though.
  • New crosses of SCM males to MTD virgins (2 vials).
  • Examine current crosses of all shRNA lines to MTDs — not looking good, food pretty dry (despite humid 30C).  Moved to RT.  Most vials from Monday’s crosses now have (a very few) pupae.   Flipped these into new vials.
  • Thursday’s crosses will need to get flipped later this week.  Should screen vials for dead embryos.


  • Ordered rubber cement from Amazon (should arrive Thursday). 
  • confirmed previously ordered primers are correct.   Best to check staining of 500 kb probe before starting new round of synthesis for en?


  • Imaging MP05 embryos.  Selected 44 positions (60GB)
  • drift after 5 sections imaging sessions are starting inside embryo epethilium instead of above embryo field of view.  cancelled and restarted.
  • Imaging MP07 embryos for width measurements

Cell labeling

  • coat coverglass with .5 mg/mL concavlin A (just add briefly and remove and add to next coverslip).  2 coverslips coated.
  • Plate with S2 cells, allow ~1hr to attach
  • Fix cells (over-fixed, ran 20 min).  4% PFA in 1/2 x PBT.
  • incubate coverglass 5′ in 2x SSCT in coplin jar
  • incubate coverglass at 60C in small hybe oven in coplin jar.  (skipped 2.5 min at 95C step, which is to be done before this).
  • Let cool to RT.  Add probe to clean, dry glass slide, gently drop slide on top.   Incubate O/N at 42C.

Data Analysis

  • Running preliminary STORM analysis on Cajal.  Using parameters optimized for last 3D HP1 dataset in S2 cells.  Currently running all analysis from image 1 (should probably push this back a few thousand for some images, though we can also do that retro-actively with STORMrender so it’s not a problem).
This entry was posted in Summaries and tagged , , . Bookmark the permalink.